### v1

# Fractional Programming for Stochastic Precoding over Generalized Fading Channels

Abstract
This paper seeks an efficient algorithm for stochastic precoding to maximize the long-term average weighted sum rates throughout a multiple-input multiple-output (MIMO) network. Unlike many existing works that assume a particular probability distribution model for fading channels (which is typically Gaussian), our approach merely relies on the first and second moments of fading channels. For the stochastic precoding problem, a naive idea is to directly apply the fractional programming (FP) method to the data rate inside the expectation; it does not work well because the auxiliary variables introduced by FP are then difficult to decide. To address the above issue, we propose using a lower bound to approximate the expectation of data rate. This lower bound stems from a nontrivial use of the matrix FP, and outperforms the existing lower bounds in that it accounts for generalized fading channels whose first and second moments are known. The resulting approximate problem can be efficiently solved in closed form in an iterative fashion. Furthermore, for large-scale MIMO, we improve the efficiency of the proposed algorithm by eliminating the large matrix inverse. Simulations show that the proposed stochastic precoding method outperforms the benchmark methods in both Gaussian and non-Gaussian fading channel cases.
Index Terms:
Stochastic precoding, multiple-input multiple-output (MIMO), first and second moments, long-term average weighted sum rates, matrix fractional programming (FP).
I Introduction
Stochastic precoding, a.k.a. robust beamforming, aims to combat the uncertainty of fading channels by maximizing the long-term average weighted sum rates throughout a multiple-input multiple-output (MIMO) network. While many existing works specialize to the Gaussian distribution model of fading channels, we do not assume any probability distribution model except that the first and second moments of fading channels are known. Our approach is based on a stochastic extension of the matrix fractional programming (FP) method [1, 2], with the data rate of each user treated as a function of random matrix ratio.
The first order of business in stochastic precoding is to model fading channels. A widely used model [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] assumes fading channels to be Gaussian, so it encompasses the classic Rayleigh and Rician fading models as special cases; [16, 17, 18] further develops it into a more sophisticated Bernoulli-Gaussian fading model. With the probability distribution of fading channels available, the expectation of data rate can be computed as an explicit function of precoders, so the standard optimization methods are applicable. However, the aforementioned works are limited to the specific fading channel models they assume. To enhance flexibility, [19, 20] no longer specify the probability distribution of fading channels $\mathbf{H}$, but only assume that the first moment $\mathbb{E}[\mathbf{H}]$ and the second moment $\mathbb{E}[\mathrm{vec}(\mathbf{H})\mathrm{vec}(\mathbf{H})^{H}]$ are known, where $\mathrm{vec}(\cdot)$ is the vectorization of a matrix, and $(\cdot)^{H}$ is the conjugate transpose, namely generalized fading channels. The present paper adopts this model too. Another fading channel model [21, 22, 23, 24] without specifying the distribution function is to assume the first moment $\mathbb{E}[\mathbf{H}]$ and a deviation bound $\delta>0$ so that $\|\mathbf{H}-\mathbb{E}[\mathbf{H}]\|_{\mathrm{F}}\leq\delta$, where $\|\cdot\|_{\mathrm{F}}$ is the Frobenius norm. The authors of [25] generalize this deviation-bound model further by replacing the Frobenius norm with other distortion metrics. Moreover, in the extreme, [26, 27, 28, 29, 30, 31] propose model-free approaches to stochastic precoding; they entail a large number of fading channel samples either for online optimization or for neural network training.
A variety of optimization objectives have been considered in the literature for the stochastic precoding problem. Many previous works [4, 6, 7, 11, 9, 10, 12, 14, 8, 32, 26, 29, 13, 27] seek to maximize the expected weighted sum of data rates throughout the network, and so does our work. For multiple-input single-output (MISO) secure transmission, [20] considers maximizing the signal-to-noise (SNR) of a legitimate user while suppressing that of an eavesdropper; the SNR is approximated as the signal-to-leakage-and-noise ratio (SLNR) in [5] to ease optimization. All these references except [20] require the prior knowledge of the probability distribution of fading channels. In contrast, this work assumes that only the first and second moments of fading channels are available, so the expectation of data rate is not even computable, which poses a major challenge to us. It is worth mentioning that the problem case of [20] is restricted to a simple network model with the single-stream transmission toward a single receiver, so its moment-based precoding scheme does not work for our case.
Aside from the above expectation-related problem formulations, a line of previous works view stochastic precoding from a robust optimization perspective. They can be further divided into two groups. The first group [22, 23, 24, 25, 21] maximizes the worst-case performance under the aforementioned bounded-deviation fading channel model. The other group [19, 16, 17, 18, 28, 31, 15, 3] maximizes the percentile performance based on the outage probability of data rate. This outage probability has been characterized for the Bernoulli-Gaussian fading model [16, 17, 18] and for the Rician fading model [15]. Some other works [28, 31] sidestep the outage probability analysis and learn stochastic precoding directly by neural network with uncertainty injection.
This paper pursues the maximization of the long-term average weighted sum rates. For this problem, [26] proposes a data-driven stochastic precoding algorithm that takes a large number of fading channel samples as training data. Each sample corresponds to a deterministic precoding subproblem—which is solvable by the weighted minimum mean square error (WMMSE) algorithm [33, 34]; the main idea of [26] is to iteratively average out the deterministic precoding solutions. As shown in [26], the iterations must converge to a stationary point of the original problem provided that the number of fading channel samples is sufficiently large. This online optimization method has evolved in some more recent works, e.g., [29] addresses the deterministic subproblems via neural networks, [27] combines the deterministic solutions in a proximal fashion, and [13] tailors the online optimization to the Gaussian case. In contrast, many other works propose the offline optimization methods. However, even when fading channels are assumed to be Gaussian, the expectation of data rate is still difficult to tackle directly, so [12] proposes a heuristic approximation of the original objective function, while [8, 32] suggest using an upper bound to approximate it. Nevertheless, since this is a maximization problem, it is more desirable to find a lower bound for approximation. Toward this end, [11] constructs a lower bound on the expectation of weighted sum rates, but the bound is difficult to compute in practice. Some computable lower bounds are proposed in [9, 10] for the single-cell network with a single data stream to each user. We remark that the above bounds, either upper or lower, are all designed for the Gaussian fading channel model. In contrast, the lower bound proposed in this paper accounts for a generic fading channel model which need not be Gaussian.
The main contributions of this paper can be recognized in the following three respects:
  • Stochastic Matrix FP: We develop the quadratic transform and the Lagrangian dual transform in [1, 2] to account for the stochastic matrix FP problems. This is motivated by the fact that the expectation of data rate amounts to $\mathbb{E}[\log|\mathbf{I}+\mathbf{A}\mathbf{B}^{-1}|]$, where $\mathbf{A}\mathbf{B}^{-1}$ is a random matrix ratio.
  • New Lower Bound: Equipped with the stochastic matrix FP, we construct a lower bound on the expected weighted sum rates, which is used to approximate the stochastic precoding problem in a tractable form. This new bound is endowed with much greater generality than the existing bounds [11, 9, 10] in that it accounts for any fading channel models with known first and second moments.
  • Precoding Algorithm: We show that the precoding variable can be efficiently optimized in closed form based on the first and second moments of fading channels. Furthermore, for the large-scale MIMO case, we improve the computational efficiency of the proposed algorithm by avoiding the large matrix inverse.
The remainder of this paper is organized as follows. Section II describes the system model and the stochastic precoding problem. Section III contains the main results: it first discusses why the conventional way of using FP does not work for stochastic precoding, and then proposes a new way based on a generalized lower bound. Section IV specifies the above FP method for the stochastic precoding problem considered in this paper. Section V discusses how to accelerate the algorithm when the transmitters have a large number of antennas. Section VI presents the simulation results. Lastly, the paper is concluded in Section VII.
Notation: Matrices and vectors are in boldface. For a matrix $\mathbf{A}$, $(\mathbf{A})_{i,j}$ is the $(i,j)$th entry, $\mathbf{A}^{c}$ is the complex conjugate, $\mathbf{A}^{H}$ is the conjugate transpose, $\lambda_{\max}(\mathbf{A})$ is the largest eigenvalue, and $\mathrm{vec}(\mathbf{A})$ is the vectorization of this matrix by stacking its columns, $\mathrm{Tr}(\mathbf{A})$ is the trace, $\mathbf{A}^{-1}$ is the inverse, $\sqrt{\mathbf{A}}$ is the symmetric square-root of matrix, $\Re\{\mathbf{A}\}$ retains the real part of each entry of the matrix, and $\mathrm{Diag}(\mathbf{A})$ extracts the diagonal entries in a vector form. For a vector $\mathbf{a}$, $\mathrm{diag}(\mathbf{a})$ is a diagonal matrix formed by $\mathbf{a}$. Let $\mathbb{C}^{m\times n}$ be the set of $m\times n$ complex matrices, $\mathbb{S}^{n\times n}_{+}$ the set of $n\times n$ positive semidefinite matrices, $\mathbb{S}^{n\times n}_{++}$ the set of $n\times n$ positive definite matrices, and $\mathbf{I}$ the identity matrix. For two matrices $\mathbf{A}$ and $\mathbf{B}$, let $\mathbf{A}\odot\mathbf{B}$ be their Hadamard product. Moreover, for a set of subscripted variables, we use the unsubscripted symbol to denote their collection, e.g., $\mathbf{V}=\{\mathbf{V}_{jk}\}$.
TABLE I: List of Main Variables
Symbol Definition
$L$ number of cells
$K$ number of users in each cell
$j,\ell$ index of cell or BS
$k,s$ index of user within each cell
$M_{t}$ number of transmit antennas at each BS
$M_{r}$ number of receive antennas at each user
$\mathbf{H}_{jk,\ell}$ block fading channel from BS $\ell$ to user $(j,k)$
$\mathbf{C}_{jk}$ first moment of $\mathbf{H}_{jk,j}$
$\mathbf{D}_{jk,\ell}$ second moment of $\mathbf{H}_{jk,\ell}$
$\mathbf{V}_{jk}$ precoding matrix for user $(j,k)$
$\mathcal{R}_{jk}$ data rate of user $(j,k)$
$\sigma^{2}$ background noise power
$\omega_{jk}$ date rate weight of user $(j,k)$
$P$ power constraint on each BS
$\mathbf{Y}$ auxiliary variable for quadratic transform
$\mathbf{\Gamma}$ auxiliary variable for Lagrangian dual transform
$\mathbf{Z}$ auxiliary variable for nonhomogeneous quadratic transform
$\mathbf{U}_{jk,\ell s}$ shorthand for $\mathbb{E}\big{[}\mathbf{H}_{jk,\ell}\mathbf{V}_{\ell s}\mathbf{V}_{\ell s}^{H}\mathbf{H}_{jk,\ell}^{H}\big{]}$
$\bm{\Lambda}_{j,\ell s}$ shorthand for $\mathbb{E}\big{[}\mathbf{H}_{\ell s,j}^{H}\mathbf{Y}_{\ell s}(\mathbf{I}+\mathbf{\Gamma}_{\ell s})\mathbf{Y}_{\ell s}^{H}\mathbf{H}_{\ell s,j}\big{]}.$
$\bm{\Xi}_{j}$ shorthand for $\sum_{(\ell,s)}\omega_{\ell s}\mathbf{\Lambda}_{j,\ell s}$
$\alpha_{j}$ largest eigenvalue of $\bm{\Xi}_{j}$
II System Model and Problem Formulation
Consider an $L$-cell downlink MIMO network, where each cell consists of one BS and $K$ user terminals. We refer to the BS of the $\ell$th cell as BS $\ell$, and the $k$th user terminal in the $j$th cell as user $(j,k)$, for $\ell,j=1,2,\ldots,L$ and $k=1,2,\ldots,K$. Assume that each BS has $M_{t}$ transmit antennas while each user terminal has $M_{r}$ receive antennas. Let $\mathbf{H}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ be the block fading channel from BS $\ell$ to user $(j,k)$. These $\mathbf{H}_{jk,\ell}$’s are modeled as i.i.d. random variables across the different blocks. We assume that the distribution function of $\mathbf{H}_{jk,\ell}$ is unknown or cannot be written analytically; only the first and second moments of random channels are available:
$\mathbf{C}_{jk}$ $=\mathbb{E}\big{[}\mathbf{H}_{jk,j}\big{]},$ (1)
$\mathbf{D}_{jk,\ell}$ $=\mathbb{E}\big{[}\mathrm{vec}(\mathbf{H}_{jk,\ell})\mathrm{vec}(\mathbf{H}_{jk,\ell})^{H}\big{]}.$ (2)
The goal of stochastic precoding is to design a set of precoding matrices $\{\mathbf{V}_{jk}\in\mathbb{C}^{M_{t}\times M_{r}}:\forall(j,k)\}$ that is “good” for all the blocks on average.
For a particular block, the data rate of user $(j,k)$ can be computed as [35]
$\mathcal{R}_{jk}=\log\big{|}\mathbf{I}+\mathbf{H}_{jk,j}\mathbf{V}_{jk}\mathbf{V}_{jk}^{H}\mathbf{H}_{jk,j}^{H}\mathbf{F}_{jk}^{-1}\big{|},$ (3)
where
$\mathbf{F}_{jk}=\sum_{(\ell,s)\neq(j,k)}\mathbf{H}_{jk,\ell}\mathbf{V}_{\ell s}\mathbf{V}_{\ell s}^{H}\mathbf{H}_{jk,\ell}^{H}+\sigma^{2}\mathbf{I},$ (4)
with $\sigma^{2}$ denoting the background noise power. We aim to maximize the weighted sum rates across many blocks; by the law of large numbers, the problem can be written as
$\underset{\mathbf{V}}{\text{maximize}}$ $\quad\sum^{L}_{j=1}\sum^{K}_{k=1}\mathbb{E}\Big{[}\omega_{jk}\mathcal{R}_{jk}\Big{]}$ (5a)
subject to $\quad\sum^{K}_{k=1}\text{Tr}(\mathbf{V}_{jk}\mathbf{V}_{jk}^{H})\leq P,\ \forall j,$ (5b)
where $\omega_{jk}\geq 0$ is the rate weight assigned to user $(j,k)$ in accordance with its priority, and $P>0$ is the power constraint on each BS. The expectation in the objective function of (5) is taken over the fading channels $\mathbf{H}_{jk,\ell}$. For ease of reference, the main variables are listed in Table I.
We provide two remarks about the stochastic precoding problem in (5). First, the problem is intractable because its objective function is unknown—the expectation cannot be computed without knowing the probability distribution of $\mathbf{H}_{jk,\ell}$. Second, the problem is also valid in the other stochastic precoding scenarios. For instance, we could have alternatively assumed that the randomness in $\mathcal{R}_{jk}$ is caused by the channel estimation error; In that case, $\mathbf{C}_{jk}$ is interpreted as the estimate of $\mathbf{H}_{jk,j}$, while $\mathbf{D}_{jk}$ is the variance of channel estimation error.
III Stochastic Matrix FP Methods
The matrix FP has been extensively used for the static precoding task (i.e., when the channels $\mathbf{H}_{jk,j}$ are all fixed and known precisely). It is natural to extend the existing matrix FP methods for the stochastic precoding case, as discussed in the first part of this section. Nevertheless, the second part of the section shows that applying the matrix FP to the matrix ratios directly inside expectation cannot lead to efficient optimization, because the auxiliary variables introduced by the matrix FP are difficult to decide. As such, we further construct a lower bound on the original objective function and use it to approximate the problem.
III-A Stochastic Extension of Matrix FP
We start by showing how the quadratic transform and the Lagrangian dual transform in [36] are extended for the stochastic case. To avoid complication in notation, and to present the results in full generality, we consider the following generic problem formulation. Let $\mathbf{A}_{n}(x\mid\theta^{+}_{n})\in\mathbb{S}^{M\times M}_{+}$ and $\mathbf{B}_{n}(x\mid\theta^{-}_{n})\in\mathbb{S}^{M\times M}_{++}$ be a pair of matrix-valued functions of $x$, for $n=1,2,\ldots,N$, where $\{\theta^{+}_{n},\theta^{-}_{n}\}$ is a set of implicit random variables that impact the function outputs. To ease notation, the arguments $(x\mid\theta^{+}_{n})$ and $(x\mid\theta^{-}_{n})$ are omitted in the rest of this section, unless we wish to emphasize their specific values. Let us consider an abstraction of the stochastic precoding problem:
$\underset{x\in\mathcal{X}}{\text{maximize}}$ $\quad\mathbb{E}\Bigg{[}\sum^{N}_{n=1}\omega_{n}\log\Big{|}\mathbf{I}+\mathbf{A}_{n}\mathbf{B}_{n}^{-1}\Big{|}\Bigg{]},$ (6)
where $\mathcal{X}$ is a nonempty constraint set and each $\omega_{n}\geq 0$ is a positive weight. Following the Lagrangian dual transform [36], we try to move the matrix ratio term out of the logdet function, as stated in the following theorem.
Theorem 1
Problem (6) is equivalent to
$\underset{x\in\mathcal{X}}{\text{maximize}}\;\mathbb{E}\bigg{[}\sup_{\bm{\Gamma}}\sum^{N}_{n=1}\omega_{n}\Big{(}\log|\mathbf{I}+\bm{\Gamma}_{n}|-\mathrm{Tr}\big{(}\bm{\Gamma}_{n}\big{)}$
$\qquad\quad+\mathrm{Tr}\big{(}(\mathbf{I}+\bm{\Gamma}_{n})\sqrt{\mathbf{A}_{n}}^{H}(\mathbf{A}_{n}+\mathbf{B}_{n})^{-1}\sqrt{\mathbf{A}_{n}}\big{)}\Big{)}\bigg{]}$ (7)
in the sense that $x$ is a global solution (or a stationary point solution) to (6) iff it is a global solution (or a stationary point solution) to (7), where a positive semidefinite auxiliary variable $\bm{\Gamma}_{n}\in\mathbb{S}^{M\times M}_{+}$ is introduced for each matrix ratio.
Proof:
For any feasible $x\in\mathcal{X}$, the optimal $\bm{\Gamma}_{n}$ in (7) is given by
$\bm{\Gamma}^{\star}_{n}=\sqrt{\mathbf{A}_{n}}^{H}\mathbf{B}_{n}^{-1}\sqrt{\mathbf{A}_{n}}.$ (8)
Substituting $\bm{\Gamma}^{\star}_{n}$ into problem (7) recovers problem (6). ∎
When $\bm{\Gamma}$ is fixed, we only need to look at the second line of (7) for optimizing $x$. This subproblem boils down to
$\underset{x\in\mathcal{X}}{\text{maximize}}$ $\quad\mathbb{E}\Bigg{[}\sum^{N}_{n=1}\mathrm{Tr}\big{(}\mathbf{Q}_{n}\sqrt{\mathbf{A}_{n}}^{H}\widetilde{\mathbf{B}}_{n}^{-1}\sqrt{\mathbf{A}_{n}}\big{)}\Bigg{]},$ (9)
where $\mathbf{Q}_{n}\in\mathbb{S}^{M\times M}_{++}$ is a positive definite matrix, and $\widetilde{\mathbf{B}}_{n}\in\mathbb{S}^{M\times M}_{++}$ is a random matrix-valued function of $x$. We now mimic the quadratic transform in [36], by decoupling each matrix ratio, as stated in the following theorem.
Theorem 2
Problem (9) is equivalent to
$\underset{x\in\mathcal{X}}{\text{maximize}}$ $\quad\mathbb{E}\bigg{[}\sup_{\mathbf{Y}}\sum^{N}_{n=1}\mathrm{Tr}\Big{(}\mathbf{Q}_{n}\big{(}\sqrt{\mathbf{A}_{n}}^{H}\mathbf{Y}_{n}$
$\qquad\qquad+\mathbf{Y}_{n}^{H}\sqrt{\mathbf{A}_{n}}-\mathbf{Y}_{n}^{H}\widetilde{\mathbf{B}}_{n}\mathbf{Y}_{n}\big{)}\Big{)}\bigg{]}$ (10)
in the sense that $x$ is a global solution (or a stationary point solution) to (9) iff it is a global solution (or a stationary point solution) to (10), where the auxiliary variable $\mathbf{Y}_{n}\in\mathbb{C}^{M\times M}$ is introduced for each matrix ratio.
Proof:
For any feasible $x\in\mathcal{X}$, the optimal $\mathbf{Y}_{n}$ in (10) is given by
$\mathbf{Y}_{n}^{\star}=\Big{(}\widetilde{\mathbf{B}}_{n}(x\mid\theta^{+}_{n},\theta^{-}_{n})\Big{)}^{-1}\sqrt{\mathbf{A}_{n}(x\mid\theta^{+}_{n})}.$ (11)
Substituting $\mathbf{Y}_{n}^{\star}$ into problem (10) recovers problem (9). ∎
III-B Proposed Lower Bound
Combining the transformations in the above two theorems, we convert problem (6) to
$\underset{x\in\mathcal{X}}{\text{maximize}}$ $\quad f(x):=\mathbb{E}\bigg{[}\sup_{\bm{\Gamma},\mathbf{Y}}g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+},\theta^{-})\bigg{]},$ (12)
where
$g(x,\bm{\Gamma},\mathbf{Y}$ $\mid\theta^{+},\theta^{-})=\sum^{N}_{n=1}\omega_{n}\Big{[}\log|\mathbf{I}+\bm{\Gamma}_{n}|-\mathrm{Tr}(\bm{\Gamma}_{n})$
$\quad+\mathrm{Tr}\big{(}(\mathbf{I}+\bm{\Gamma}_{n})(\sqrt{\mathbf{A}_{n}}^{H}\mathbf{Y}_{n}+\mathbf{Y}_{n}^{H}\sqrt{\mathbf{A}_{n}})\big{)}$
$\quad-\mathrm{Tr}\big{(}(\mathbf{I}+\bm{\Gamma}_{n})\mathbf{Y}_{n}^{H}(\mathbf{A}_{n}+\mathbf{B}_{n})\mathbf{Y}_{n}\big{)}\Big{]}$ (13)
with the auxiliary variables $\mathbf{Y}_{n}\in\mathbb{C}^{M\times M}$ and $\bm{\Gamma}_{n}\in\mathbb{S}^{M\times M}_{+}$.
It is then tempting to optimize $(x,\bm{\Gamma},\mathbf{Y})$ in the above new problem in an iterative fashion, just as in the static precoding problem case. But we point out a subtle issue with this method in the following. When $x$ and $\bm{\Gamma}$ are fixed, each $\mathbf{Y}_{n}$ is optimally determined as
$\mathbf{Y}_{n}^{\star}=\Big{(}\mathbf{A}_{n}(x\mid\theta^{+}_{n})+\mathbf{B}_{n}(x\mid\theta^{-}_{n})\Big{)}^{-1}\sqrt{\mathbf{A}_{n}(x\mid\theta^{+}_{n})}.$ (14)
When $x$ and $\mathbf{Y}$ are fixed, each $\bm{\Gamma}_{n}$ is optimally determined as
$\bm{\Gamma}^{\star}_{n}=\sqrt{\mathbf{A}_{n}(x\mid\theta^{+}_{n})}^{H}\Big{(}\mathbf{B}_{n}(x\mid\theta^{-}_{n})\Big{)}^{-1}\sqrt{\mathbf{A}_{n}(x\mid\theta^{+}_{n})}.$ (15)
Note that the optimal updates of $\mathbf{Y}_{n}$ and $\bm{\Gamma}_{n}$ depend on the current realizations of the random variables $\{\theta^{+}_{n},\theta^{-}_{n}\}$, which are not available in our problem case. As a consequence, the conventional way of using FP does not work for stochastic precoding.
Like the precoding variable, the auxiliary variables $\mathbf{Y}$ and $\bm{\Gamma}$ ought to depend on the channel statistics rather than a particular realization. To remedy the problem, we propose interchanging $\mathbb{E}[\cdot]$ and $\sup_{\bm{\Gamma},\mathbf{Y}}\{\cdot\}$ in (12); subsequently, $\sup_{\bm{\Gamma},\mathbf{Y}}$ can be further incorporated into “maximize” as
$\underset{x\in\mathcal{X},\mathbf{\Gamma},\mathbf{Y}}{\text{maximize}}$ $\quad\hat{f}(x,\bm{\Gamma},\mathbf{Y}):=\mathbb{E}\bigg{[}g(x,\mathbf{Y},\bm{\Gamma}\mid\theta^{+}_{n},\theta^{-}_{n})\bigg{]},$ (16a)
subject to $\quad\bm{\Gamma}_{n}\in\mathbb{S}^{M\times M}_{+},\;\forall n$ (16b)
$\quad\mathbf{Y}_{n}\in\mathbb{C}^{M\times M},\;\forall n.$ (16c)
It can be seen that the optimal updates of $\mathbf{Y}_{n}$ and $\bm{\Gamma}_{n}$ are now independent of the realizations of $\{\theta^{+}_{n},\theta^{-}_{n}\}$.
We now justify the above problem transformation by showing that the new problem is in essence to maximize a lower bound of the original objective function in (6).
Theorem 3
For any feasible tuple $(x,\bm{\Gamma},\mathbf{Y})$, we have
$\hat{f}(x,\bm{\Gamma},\mathbf{Y})\leq\mathbb{E}\Bigg{[}\sum^{N}_{n=1}\omega_{n}\log\Big{|}\mathbf{I}+\mathbf{A}_{n}\mathbf{B}_{n}^{-1}\Big{|}\Bigg{]}.$ (17)
Proof:
Because of the equivalence between (6) and (12), it suffices to show that
$\hat{f}(x,\bm{\Gamma},\mathbf{Y})\leq f(x).$ (18)
We then have
$\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ $\leq\sup_{\bm{\Gamma},\mathbf{Y}}\hat{f}(x,\bm{\Gamma},\mathbf{Y})$ (19)
$=\sup_{\bm{\Gamma},\mathbf{Y}}\mathbb{E}[g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+}_{n},\theta^{-}_{n})]$ (20)
$\leq\mathbb{E}\bigg{[}\sup_{\bm{\Gamma},\mathbf{Y}}g(x,\bm{\Gamma},\mathbf{Y}\mid\theta^{+}_{n},\theta^{-}_{n})\bigg{]}$ (21)
$=f(x).$ (22)
The proof is complete. ∎
In the next section, we show that the new problem (16) allows the precoding variable and the auxiliary variables to be efficiently optimized based on the first and second moments of fading channels.
IV FP-Based Stochastic Precoding
$\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})=\\ \sum_{(j,k)}\omega_{jk}\Bigg{[}\log\big{|}\mathbf{I}+\mathbf{\Gamma}_{jk}\big{|}-\mathrm{Tr}\big{(}\mathbf{\Gamma}_{jk}\big{)}+\mathrm{Tr}\Bigg{(}\big{(}\mathbf{I}+\mathbf{\Gamma}_{jk}\big{)}\Bigg{(}2\Re\big{\{}\mathbf{V}_{jk}^{H}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}\big{\}}-\mathbf{Y}_{jk}^{H}\Bigg{(}\sum_{(\ell,s)}\mathbf{U}_{jk,\ell s}+\sigma^{2}\mathbf{I}\Bigg{)}\mathbf{Y}_{jk}\Bigg{)}\Bigg{)}\Bigg{]}$ (23)
We now specialize the above matrix FP method for the stochastic precoding problem (5). In light of the following correspondences:
$\sqrt{\mathbf{A}_{n}(x\mid\theta^{+}_{n})}$ $\;\Longleftrightarrow\;\mathbf{H}_{jk,j}\mathbf{V}_{jk},$ (24)
$\mathbf{B}_{n}(x\mid\theta^{-}_{n})$ $\;\Longleftrightarrow\;\mathbf{F}_{jk},$ (25)
problem (5) is converted to
$\underset{\mathbf{V},\bm{\Gamma},\mathbf{Y}}{\text{maximize}}$ $\quad\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ (26a)
subject to $\quad\sum^{K}_{k=1}\text{Tr}(\mathbf{V}_{jk}\mathbf{V}_{jk}^{H})\leq P,\;\forall j$ (26b)
$\quad\,\bm{\Gamma}_{jk}\in\mathbb{S}^{M_{r}\times M_{r}}_{+},\;\forall(j,k)$ (26c)
$\quad\,\mathbf{Y}_{jk}\in\mathbb{C}^{M_{r}\times M_{r}},\;\forall(j,k),$ (26d)
where the new objective function $\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ is shown in (23) at the top of next page, with the shorthand
$\mathbf{U}_{jk,\ell s}=\mathbb{E}\big{[}\mathbf{H}_{jk,\ell}\mathbf{V}_{\ell s}\mathbf{V}_{\ell s}^{H}\mathbf{H}_{jk,\ell}^{H}\big{]}.$ (27)
Each entry of $\mathbf{U}_{jk,\ell s}\in\mathbb{S}^{M_{r}\times M_{r}}_{+}$ can be obtained from the current $\mathbf{V}$ and the second moment $\mathbf{D}_{jk,\ell}$ as
$\big{(}\mathbf{U}_{jk,\ell s}\big{)}_{m,n}$
$=\sum_{m^{\prime}=1}^{M_{t}}\sum_{n^{\prime}=1}^{M_{t}}\big{(}\mathbf{V}_{\ell s}\mathbf{V}^{H}_{\ell s}\big{)}_{m^{\prime},n^{\prime}}\mathbb{E}\big{[}\big{(}\mathbf{H}_{jk,\ell}\big{)}_{m,m^{\prime}}\big{(}\mathbf{H}^{c}_{jk,\ell}\big{)}_{n,n^{\prime}}\big{]}$
$=\sum_{m^{\prime}=1}^{M_{t}}\sum_{n^{\prime}=1}^{M_{t}}\Big{(}\big{(}\mathbf{V}_{\ell s}\mathbf{V}^{H}_{\ell s}\big{)}_{m^{\prime},n^{\prime}}$
×$\qquad\times(\mathbf{D}_{jk,\ell})_{(m^{\prime}-1)M_{r}+m,(n^{\prime}-1)M_{r}+n}\Big{)}$ (28)
We then optimize $(\mathbf{V},\bm{\Gamma},\mathbf{Y})$ in an iterative fashion. When $\mathbf{V}$ and $\mathbf{Y}$ are both held fixed, each $\bm{\Gamma}_{jk}$ can be optimally determined as
$\mathbf{\Gamma}_{jk}^{\star}=\mathbf{V}_{jk}^{H}\mathbf{C}_{jk}^{H}\Bigg{(}\sigma^{2}\mathbf{I}+\sum_{(\ell,s)}\mathbf{U}_{jk,\ell s}\\ -\mathbf{C}_{jk}\mathbf{V}_{jk}\mathbf{V}_{jk}^{H}\mathbf{C}_{jk}^{H}\Bigg{)}^{-1}\mathbf{C}_{jk}\mathbf{V}_{jk}.$ (29)
When $\mathbf{V}$ and $\bm{\Gamma}$ are both held fixed, each $\mathbf{Y}_{jk}$ can be optimally determined as
$\mathbf{Y}_{jk}^{\star}=\Bigg{(}\sigma^{2}\mathbf{I}+\sum_{(\ell,s)}\mathbf{U}_{jk,\ell s}\Bigg{)}^{-1}\mathbf{C}_{jk}\mathbf{V}_{jk}.$ (30)
After $\bm{\Gamma}$ has been optimized according to (29) for the current $\mathbf{Y}$, the solution of each $\mathbf{V}_{jk}$ is given by
$\mathbf{V}_{jk}=\Bigg{(}\eta_{j}\mathbf{I}+\sum_{(\ell,s)}\omega_{\ell s}\bm{\Lambda}_{j,\ell s}\Bigg{)}^{-1}\omega_{jk}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}(\mathbf{I}+\mathbf{\Gamma}_{jk}),$ (31)
where the Lagrangian multiplier $\eta_{j}\geq 0$ for the power constraint is optimally determined as
$\eta_{j}^{\star}=\min\Bigg{\{}\eta_{j}\geq 0:\sum^{K}_{k=1}\mathrm{Tr}(\mathbf{V}_{jk}\mathbf{V}_{jk}^{H})\leq P\Bigg{\}},$ (32)
and the matrix $\bm{\Lambda}_{j,\ell s}\in\mathbb{S}^{M_{t}\times M_{t}}_{+}$ is given by
$\bm{\Lambda}_{j,\ell s}=\mathbb{E}\big{[}\mathbf{H}_{\ell s,j}^{H}\mathbf{Y}_{\ell s}(\mathbf{I}+\mathbf{\Gamma}_{\ell s})\mathbf{Y}_{\ell s}^{H}\mathbf{H}_{\ell s,j}\big{]}.$ (33)
We can use $\mathbf{D}_{jk,\ell}$ to recover each entry of $\bm{\Lambda}_{j,\ell s}$ as
$\big{(}\bm{\Lambda}_{j,\ell s}\big{)}_{m,n}$
$=\sum_{m^{\prime}=1}^{M_{r}}\sum_{n^{\prime}=1}^{M_{r}}\Big{(}\big{(}\mathbf{Y}_{\ell s}(\mathbf{I}+\mathbf{\Gamma}_{\ell s})\mathbf{Y}_{\ell s}^{H}\big{)}_{m^{\prime},n^{\prime}}$
×$\qquad\qquad\qquad\quad\times\mathbb{E}\big{[}\big{(}\mathbf{H}^{c}_{\ell s,j}\big{)}_{m^{\prime},m}\big{(}\mathbf{H}_{\ell s,j}\big{)}_{n^{\prime},n}\big{]}\Big{)}$
$=\sum_{m^{\prime}=1}^{M_{r}}\sum_{n^{\prime}=1}^{M_{r}}\Big{(}\big{(}\mathbf{Y}_{\ell s}(\mathbf{I}+\mathbf{\Gamma}_{\ell s})\mathbf{Y}_{\ell s}^{H}\big{)}_{m^{\prime},n^{\prime}}$
×$\qquad\qquad\qquad\times(\mathbf{D}_{\ell s,j}^{c})_{(m-1)M_{r}+m^{\prime},(n-1)M_{r}+n^{\prime}}\Big{)}.$ (34)
Algorithm 1 summarizes the above steps for stochastic precoding. Its per-iteration complexity equals $\mathcal{O}(L^{2}K^{2}M_{r}^{2}M_{t}^{2}+LKM_{r}^{3}+LM_{t}^{3})$. The following theorem provides a convergence analysis for this algorithm.
Algorithm 1 Stochastic Precoding Based on Matrix FP
1:  input: first and second moments $\{\mathbf{C},\mathbf{D}\}$ of fading channels
2:  repeat
3:     compute $\mathbf{U}$ by (28) and then update $\mathbf{Y}$ by (30)
4:     update $\bm{\Gamma}$ by (29)
5:     compute $\bm{\Lambda}$ by (34) and then update $\mathbf{V}$ by (31)
6:  until the objective value converges
7:  output: precoding matrix $\mathbf{V}$
$\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})=\sum_{(j,k)}\Bigg{(}\omega_{jk}\log\big{|}\mathbf{I}+\mathbf{\Gamma}_{jk}\big{|}-\omega_{jk}\mathrm{Tr}\big{(}\mathbf{\Gamma}_{jk}\big{)}\\ +\mathrm{Tr}\Bigg{(}2\omega_{jk}\big{(}\mathbf{I}+\mathbf{\Gamma}_{jk}\big{)}\Re\big{\{}\mathbf{V}_{jk}^{H}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}\big{\}}-\omega_{jk}\sigma^{2}\big{(}\mathbf{I}+\mathbf{\Gamma}_{jk}\big{)}\mathbf{Y}_{jk}^{H}\mathbf{Y}_{jk}-\mathbf{V}_{jk}^{H}\Bigg{(}\sum_{(\ell,s)}\omega_{\ell s}\mathbf{\Lambda}_{j,\ell s}\Bigg{)}\mathbf{V}_{jk}\Bigg{)}\Bigg{)}$ (35)
$\zeta(\mathbf{V},\mathbf{\Gamma},\mathbf{Y},\mathbf{Z})=\sum_{(j,k)}\bigg{(}\omega_{jk}\log|\mathbf{I}+\mathbf{\Gamma}_{jk}|-\omega_{jk}\mathrm{Tr}(\mathbf{\Gamma}_{jk})+2\mathrm{Tr}\Big{(}\Re\Big{\{}\omega_{jk}(\mathbf{I}+\mathbf{\Gamma}_{jk})\mathbf{V}_{jk}^{H}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}+\mathbf{V}_{jk}^{H}(\alpha_{j}\mathbf{I}-\bm{\Xi}_{j})\mathbf{Z}_{jk}\Big{\}}\Big{)}\\ +\mathrm{Tr}\Big{(}\mathbf{Z}_{jk}^{H}(\bm{\Xi}_{j}-\alpha_{j}\mathbf{I})\mathbf{Z}_{jk}-\alpha_{j}\mathbf{V}_{jk}^{H}\mathbf{V}_{jk}-\omega_{jk}\sigma^{2}(\mathbf{I}+\mathbf{\Gamma}_{jk})\mathbf{Y}_{jk}^{H}\mathbf{Y}_{jk}\Big{)}\bigg{)}$ (36)
Proposition 1
Algorithm 1 guarantees convergence to a stationary point $(\mathbf{V},\bm{\Gamma},\mathbf{Y})$ of problem (26). Moreover, the new objective in (26) is a lower bound to the original objective:
$\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})\leq\sum^{L}_{j=1}\sum^{K}_{k=1}\mathbb{E}\Big{[}\omega_{jk}\mathcal{R}_{jk}\Big{]}.$ (37)
Proof:
The convergence result follows by the property of FP as shown in [37]. The lower bound result follows by Theorem 3. ∎
Remark 1
The computations of $\mathbf{U}$ and $\bm{\Lambda}$ in Algorithm 1 can be further simplified when fading channels are known to be Rayleigh or Rician. In that case, each $\mathbf{H}_{jk,\ell}$ has a Gaussian distribution:
$\mathbf{H}_{jk,\ell}=\widetilde{\mathbf{H}}_{jk,\ell}+\mathbf{W}_{jk,\ell}\odot\mathbf{X}_{jk,\ell},$ (38)
where $\widetilde{\mathbf{H}}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ and $\mathbf{W}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ are fixed matrices, and $\mathbf{X}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ is a random matrix with each entry drawn i.i.d. from the standard complex Gaussian distribution $\mathcal{CN}(0,1)$. Then $\mathbf{U}$ and $\bm{\Lambda}$ are directly given by
$\mathbf{U}_{jk,\ell s}=\widetilde{\mathbf{H}}_{jk,\ell}\mathbf{V}_{\ell s}\mathbf{V}_{\ell s}^{H}\widetilde{\mathbf{H}}_{jk,\ell}^{H}$
$\;+\mathrm{diag}\Big{(}\big{(}\mathbf{W}_{jk,\ell}\odot\mathbf{W}_{jk,\ell}\big{)}\mathrm{Diag}\big{(}\mathbf{V}_{\ell s}\mathbf{V}_{\ell s}^{H}\big{)}\Big{)},$ (39)
$\bm{\Lambda}_{j,\ell s}=\widetilde{\mathbf{H}}_{jk,\ell}^{H}\mathbf{Y}_{\ell s}(\mathbf{I}+\mathbf{\Gamma}_{\ell s})\mathbf{Y}_{\ell s}^{H}\widetilde{\mathbf{H}}_{jk,\ell}$
$\;+\mathrm{diag}\Big{(}\big{(}\mathbf{W}_{jk,\ell}^{H}\odot\mathbf{W}_{jk,\ell}^{H}\big{)}\mathrm{Diag}\big{(}\mathbf{Y}_{\ell s}(\mathbf{I}+\mathbf{\Gamma}_{\ell s})\mathbf{Y}_{\ell s}^{H}\big{)}\Big{)},$ (40)
so we no longer need to compute these matrices entry by entry as in (28) and (34). As a result, the per-iteration complexity of Algorithm 1 reduces to $\mathcal{O}(L^{2}K^{2}M_{r}M_{t}^{2}+L^{2}K^{2}M_{r}^{2}M_{t}+LKM_{r}^{3}+LM_{t}^{3})$. In particular, if we further let $M_{r}=1$ then Algorithm 1 reduces to the precoding method in [9].
V Acceleration for Large-Scale MIMO
In this section we cope with such a challenge: for the large-scale MIMO case with $M_{t}$ being a large number, the $M_{t}\times M_{t}$ matrix inverse in (31) renders the update of $\mathbf{V}$ computationally formidable in Step 5 of Algorithm 1. The main idea is to get rid of this $M_{t}\times M_{t}$ matrix inverse by further modifying the matrix FP. We start by introducing a useful inequality from [38, 39].
Lemma 1
Consider any two ${M\times M}$ Hermitian matrices $\mathbf{L},\mathbf{K}$ satisfying $\mathbf{L}\preceq\mathbf{K}$. For any $\mathbf{X},\mathbf{Z}\in\mathbb{C}^{M\times N}$, we have
$\mathrm{Tr}\big{(}\mathbf{X}^{H}\mathbf{L}\mathbf{X}\big{)}\leq\mathrm{Tr}\big{(}\mathbf{X}^{H}\mathbf{K}\mathbf{X}+2\Re\{\mathbf{X}^{H}(\mathbf{L}-\mathbf{K})\mathbf{Z}\}\\ +\mathbf{Z}^{H}(\mathbf{K}-\mathbf{L})\mathbf{Z}\big{)},$ (41)
where the equality holds if $\mathbf{Z}=\mathbf{X}$.
To see how the above inequality may assist with our problem case, we first rewrite $\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ in (23) in a quadratic form of $\mathbf{V}_{jk}$, as shown in (35). Now observe that the $M_{t}\times M_{t}$ matrix inverse in (31) is caused by the quadratic term $\mathbf{V}_{jk}^{H}(\sum_{(\ell,s)}\omega_{\ell s}\mathbf{\Lambda}_{j,\ell s})\mathbf{V}_{jk}$ in (35). Let us treat $\sum_{(\ell,s)}\omega_{\ell s}\mathbf{\Lambda}_{j,\ell s}$ as $\mathbf{L}$ in (41). The main idea here is to replace this matrix with a “simpler” matrix $\mathbf{K}$ in the quadratic term by using Lemma 1; since we seek to eliminate matrix inverse, it is desirable to let $\mathbf{K}=\alpha\mathbf{I}$ so that its inverse is directly given by $\mathbf{K}^{-1}=(1/\alpha)\mathbf{I}$. But how do we choose the factor $\alpha$ to meet the condition $\mathbf{L}\preceq\mathbf{K}$ in Lemma 1? As suggested in [40], we may set $\alpha$ to the largest eigenvalue of $\mathbf{L}$.
Algorithm 2 Fast Stochastic Precoding for Large-Scale MIMO
1:  input: first and second moments $\{\mathbf{C},\mathbf{D}\}$ of fading channels
2:  repeat
3:     compute $\mathbf{U}$ by (28) and then update $\mathbf{Y}$ by (30)
4:     update $\bm{\Gamma}$ by (29)
5:     update $\mathbf{Z}$ by (46)
6:     compute $\bm{\Lambda}$ by (34) and then update $\mathbf{V}$ by (51)
7:  until the objective value converges
8:  output: precoding matrix $\mathbf{V}$
By applying the inequality (41) to $\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ in (23), we obtain a lower bound
$\zeta(\mathbf{V},\mathbf{\Gamma},\mathbf{Y},\mathbf{Z})\leq\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y}),$ (41)
which is shown in (36) with the shorthands
$\bm{\Xi}_{j}=\sum_{(\ell,s)}\omega_{\ell s}\mathbf{\Lambda}_{j,\ell s}$ (43)
and
$\alpha_{j}=\lambda_{\max}(\bm{\Xi}_{j}).$ (44)
Problem (26) is now turned to
$\underset{\mathbf{V},\mathbf{Z},\bm{\Gamma},\mathbf{Y}}{\text{maximize}}$ $\quad\zeta(\mathbf{V},\mathbf{Z},\mathbf{\Gamma},\mathbf{Y})$ (45a)
[…]
$\quad\,\mathbf{Z}_{jk}\in\mathbb{C}^{M_{t}\times M_{r}},\;\forall(j,k).$ (45c)
As before, we optimize these variables in an iterative fashion. When $(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ are fixed, as shown in Lemma 1, each $\mathbf{Z}_{j}$ can be optimally determined as
$\mathbf{Z}_{jk}^{\star}=\mathbf{V}_{jk},$ (46)
The iterative optimal updates of $\mathbf{Y}$ and $\bm{\Gamma}$ are still given by (30) and (29), respectively. Moreover, for fixed $(\mathbf{Y},\mathbf{\Gamma},\mathbf{Z})$, after removing the constant terms, problem (45) can be divided into a set of subproblems on a per-cell basis:
$\underset{\mathbf{V}_{j1},\dots,\mathbf{V}_{jK}}{\text{minimize}}$ $\quad\sum^{K}_{k=1}\mathrm{Tr}\Big{(}2\Re\Big{\{}\omega_{jk}(\mathbf{I}+\mathbf{\Gamma}_{jk})\mathbf{V}_{jk}^{H}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}$
$\quad+\mathbf{V}_{jk}^{H}(\alpha_{j}\mathbf{I}-\bm{\Xi}_{j})\mathbf{Z}_{jk}\Big{\}}-\alpha_{j}\mathbf{V}_{jk}^{H}\mathbf{V}_{jk}\Big{)}$ (47a)
subject to $\quad\sum^{K}_{k=1}\mathrm{Tr}(\mathbf{V}_{jk}\mathbf{V}_{jk}^{H})\leq P,$ (47b)
for $j=1,2,\ldots,L$. The optimal solution of $\mathbf{V}_{jk}$ is given by
$\mathbf{V}_{jk}^{\star}=\frac{1}{\alpha_{j}+\eta_{j}^{\star}}\bigg{(}\omega_{jk}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}(\mathbf{I}+\mathbf{\Gamma}_{jk})+\big{(}\alpha_{j}\mathbf{I}-\mathbf{\Xi}_{j}\big{)}\mathbf{Z}_{jk}\bigg{)},$ (48)
where the optimal Lagrange multiplier $\eta_{j}^{\star}$ is still computed as in (32). After a bit algebra, it can be shown that the above solution boils down to
$\mathbf{V}^{\star}_{jk}=\left\{\begin{array}[]{ll}{\bm{\Psi}}_{jk}&\text{if}\;\sum^{K}_{k=1}\mathrm{Tr}(\bm{\Psi}_{jk}^{H}\bm{\Psi}_{jk})\leq P\vspace{0.5em}\\ \beta_{j}\bm{\Psi}_{jk}&\text{otherwise},\end{array}\right.$ (51)
where
$\bm{\Psi}_{jk}$ $=\mathbf{Z}_{jk}+\frac{1}{\alpha_{j}}\Big{(}\omega_{jk}\mathbf{C}_{jk}^{H}\mathbf{Y}_{jk}(\mathbf{I}+\mathbf{\Gamma}_{jk})-\mathbf{\Xi}_{j}\mathbf{Z}_{jk}\Big{)}$ (52)
and
$\beta_{j}=\sqrt{\frac{P}{\sum^{K}_{k=1}\mathrm{Tr}(\bm{\Psi}_{jk}^{H}\bm{\Psi}_{jk})}}.$ (53)
Algorithm 2 summarizes the above steps. Its per-iteration complexity equals $\mathcal{O}(L^{2}K^{2}M_{r}^{2}M_{t}^{2}+LKM_{r}^{3})$.
The following proposition shows that Algorithm 2 converges to a stationary point solution not only of problem (47) but also of problem (26), so the problem approximation based on Lemma 1 does not cause performance loss.
Proposition 2
Algorithm 2 converges to $(\mathbf{V}_{0},\bm{\Gamma}_{0},\mathbf{Y}_{0},\mathbf{Z}_{0})$ which is a stationary point of problem (47). Moreover, $(\mathbf{V}_{0},\bm{\Gamma}_{0},\mathbf{Y}_{0})$ is a stationary point of problem (26).
Proof:
We only sketch the basic ideas which follow the existing work [40] closely. First, Algorithm 2 can be recognized as a minorization-maximization (MM) method for problem (47); because each step of Algorithm 2 boils down to solving a convex subproblem, the convergence to a stationary point can be verified by the MM theory [40]. Moreover, using the inequality in Lemma 1 to bound $\hat{f}(\mathbf{V},\mathbf{\Gamma},\mathbf{Y})$ can be thought of as constructing a surrogate function for problem (26), so each stationary point of problem (47) can be mapped to a stationary point of problem (26) according to the MM theory. ∎
VI Simulations
Figure 1: Lower bounds v.s. original objective for different noise power levels.
Figure 2: Lower bounds v.s. original objective for different values of the temporal correlation coefficient $\rho_{jk}$.
VI-A Settings
We consider two fading models: the Rayleigh fading which is Gaussian, and the Nakagami-$m$ fading which is non-Gaussian. We first describe the Rayleigh fading model from [11, 10, 13]. The channel is modeled as
$\mathbf{H}_{jk,\ell}=\rho_{jk}\bar{\mathbf{H}}_{jk,\ell}+\sqrt{1-(\rho_{jk})^{2}}\mathbf{W}_{jk,\ell}\odot\mathbf{X}_{jk,\ell},$ (54)
where $\bar{\mathbf{H}}_{jk,\ell}$ is the static component, $\mathbf{W}_{jk,\ell}\odot\mathbf{X}_{jk,\ell}$ is the random component, and $\rho_{jk}\in(0,1]$ is the tradeoff factor. These parameters are further specified in the following. The real matrix $\mathbf{W}_{jk,\ell}\in\mathbb{R}^{M_{r}\times M_{t}}$ reflects on the large-scale fading, each entry of which is computed as $10^{-\xi/10}$ where $\xi=0.5\times(128.1+37.6\log_{10}d+\tau)$, $d$ denoting the distance in km between the corresponding pair of BS and user, $\tau$ being a zero-mean Gaussian random variable with a standard deviation of 8 dB for the shadowing effect. The fixed matrix $\bar{\mathbf{H}}_{jk,\ell}$ is a realization of $\mathbf{W}_{jk,\ell}\odot\mathbf{X}_{jk,\ell}$ at the beginning of the test. The random complex matrix $\mathbf{X}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$ has each entry independently drawn from the standard complex Gaussian distribution $\mathcal{CN}(0,1)$. The tradeoff factor $\rho_{jk}$, a.k.a. the temporal correlation coefficient, is generated by Jake’s model [41]. We consider $1000$ independent blocks and evaluate their average rates.
We also consider the Nakagami-$m$ fading model as in [42]:
$\mathbf{H}_{jk,\ell}=\rho_{jk}\bar{\mathbf{H}}_{jk,\ell}+\sqrt{1-(\rho_{jk})^{2}}\mathbf{M}_{jk,\ell}.$ (55)
Now the random component becomes a Nakagami-$m$ random matrix $\mathbf{M}_{jk,\ell}\in\mathbb{C}^{M_{r}\times M_{t}}$. For the $(m,n)$th entry of $\mathbf{M}_{jk,\ell}$, its phase is uniform on $(0,2\pi]$, and its magnitude is generated independently according to the Nakagami-$m$ distribution $\mathrm{Nakagami}(m,\Omega_{mn})$ [42], where the shape parameter $m=0.5$ and the mean $\Omega_{mn}$ equals the square of the large-scale fading $\xi$ as defined earlier for the Rayleigh fading model. Again, $\bar{\mathbf{H}}_{jk,\ell}$ is a realization of $\mathbf{M}_{jk,\ell}$ and is fixed ever since. We still average out the performance across $1000$ independent blocks.
(a) Single-cell Rayleigh fading case
(b) Multi-cell Rayleigh fading case
(c) Single-cell Nakagami-$m$ fading case
(d) Multi-cell Nakagami-$m$ fading case
Figure 3: The long-term average sum rates achieved by the different algorithms under the different fading models and under the different noise power levels.
We further specify the network setting. The radius of each cell equals $300$ meters. We let $P=30$ dBm and $\omega_{jk}=1$. The noise power is $-90$ dBm by default. For the multi-cell case, we consider 7 wrapped-around cells. We let $M_{t}=64$, $K=32$, and $M_{r}=2$ by default for the single-cell case, and $M_{t}=32$, $K=16$, and $M_{r}=2$ by default for the multi-cell case. Moreover, the benchmark methods for stochastic precoding include:
  • WMMSE: Run the WMMSE algorithm [34] based on the static components of fading channels.
  • Robust Linear Precoding Design (RLPD): Maximize an approximation of the original objective function [11].
  • Stochastic WMMSE (SWMMSE): A model-free precoding method in [26] based on fading channel samples.
VI-B Test Results
We begin by evaluating the tightness of the proposed lower bound in (3). Fig. 1 and Fig. 2 consider the Rayleigh fading model. We remark that the lower bound in [11] is not included in the above two figures because it is uncomputable. It can be seen that the proposed bound is much tighter than the existing one in [12].
Next, we compare the expected sum rates achieved by different algorithms under the different values of noise power $\sigma^{2}$ in Fig. 3. Note that RLPD is limited to the single-cell Rayleigh fading case, so it only appears in Fig. 3(a). For SWMMSE, we provide a total of $1000$ channel samples for online optimization. Because Algorithm 1 and Algorithm 2 only differ in the computational complexity, they produce the same performance curves in Fig. 3. Observe that the WMMSE algorithm that only depends on the static components of fading channels lead to the worst performance. The proposed method attains much better performance, which is about 30% higher than that of WMMSE. Observe also that SWMMSE yields fairly nice performance in the single-cell case, which is almost equally good as the proposed method. However, SWMMSE becomes much worse in the presence of multiple cells since then it requires many more channel samples. In practice, online optimization by SWMMSE can be quite costly. RLPD has good performance, but unfortunately it is limited to the single-cell Rayleigh fading case.
(a) Single-cell Rayleigh fading case
(b) Multi-cell Rayleigh fading case
(c) Single-cell Nakagami-$m$ fading case
(d) Multi-cell Nakagami-$m$ fading case
Figure 4: The convergence of the different algorithms in terms of the number of iterations under the different fading models.
(a) Single-cell Rayleigh fading case
(b) Multi-cell Rayleigh fading case
(c) Single-cell Nakagami-$m$ fading case
(d) Multi-cell Nakagami-$m$ fading case
Figure 5: The convergence of the different algorithms in terms of elapsed time under the different fading models.
Moreover, we compare the speed of convergence for the different algorithms. We now draw Algorithm 1 and Algorithm 2 separately. First, we evaluate the speed in iterations. As shown in Fig. 4, WMMSE cannot even yield monotonic improvement for the objective value. It can be seen that SWMMSE converges slowly; this low efficiency can be expected since online optimization mines for solution from massive channel samples. In particular, note that Algorithm 2 is even slower than Algorithm 1. Actually, this is not a surprising result. When we talk about acceleration in Section V, we mean the simplification of each iteration by eliminating matrix inverse. However, as shown in [40], this simplification is at the cost of slowing down the convergence in iterations. (As shown in [40], this drawback can be remedied by incorporating momentum into the FP method, but this is beyond the scope of this work.) Nevertheless, the acceleration in computation outweighs the deceleration in iterations. In plain words, even though it now takes more iterations to converge, the run time of each iteration becomes much shorter, and hence the overall time consumption is still reduced significantly by Algorithm 2. The results from Fig. 5 agree with the above discussion. Observe that Algorithm 2 is about 3 times faster than Algorithm 1. Actually, when Algorithm 1 just finishes one iteration, Algorithm 2 already reaches convergence. We further compare their run times in Table II by using different numbers of antennas. Observe that the efficiency advantage of Algorithm 2 over Algorithm 1 becomes larger when the number of antennas increases, so Algorithm 2 is well suited for the large-scale MIMO network.
TABLE II: Per-Iteration Run Time of Different Methods.
Average Running Time (second)
Algorithm $M_{t}=32$ $M_{t}=64$ $M_{t}=128$ $M_{t}=256$
WMMSE 0.0960 0.2735 1.1187 5.4836
Algorithm 1 0.1209 0.3707 1.4573 6.8598
Algorithm 2 0.0039 0.0064 0.0167 0.1597
RLPD 0.0656 0.1555 0.4953 3.0894
SWMMSE 0.1199 0.3974 1.4926 5.0732
VII Conclusion
This work investigates the stochastic precoding design by using the first and second moments of fading channels. As compared to the existing works that assume a particular channel distribution, our method has the advantage of working for a generic fading model; as compared to the existing works that adopt a model-free approach, our method has the advantage of providing more efficient optimization. The heart of our method lies in a stochastic extension of the existing matrix FP methods. Although this paper focuses on the downlink network, the proposed algorithm can be readily adapted to other network topologies such as device-to-device, uplink, and cell-free.